# GaussianBeV : 3D Gaussian Representation meets Perception Models for BeV Segmentation

### v1

@@ Abstract
The Bird’s-eye View (BeV) representation is widely used for 3D perception from multi-view camera images. It allows to merge features from different cameras into a common space, providing a unified representation of the 3D scene. The key component is the view transformer, which transforms image views into the BeV. However, actual view transformer methods based on geometry or cross-attention do not provide a sufficiently detailed representation of the scene, as they use a sub-sampling of the 3D space that is non-optimal for modeling the fine structures of the environment. In this paper, we propose GaussianBeV, a novel method for transforming image features to BeV by finely representing the scene using a set of 3D gaussians located and oriented in 3D space. This representation is then splattered to produce the BeV feature map by adapting recent advances in 3D representation rendering based on gaussian splatting [12]. GaussianBeV is the first approach to use this 3D gaussian modeling and 3D scene rendering process online, i.e. without optimizing it on a specific scene and directly integrated into a single stage model for BeV scene understanding. Experiments show that the proposed representation is highly effective and place GaussianBeV as the new state-of-the-art on the BeV semantic segmentation task on the nuScenes dataset [2].
@@ 1 Introduction
In this article, we propose a novel view transformation method called GaussianBeV, which enables fine 3D modeling of scenes. Drawing on recent advances in explicit 3D representation for rendering novel views based on Gaussian Splatting (GS) [12], our method represents a scene using a set of 3D gaussians, each parameterized by a center, scale, rotation, opacity and semantic features (instead of colors in GS). Furthermore, unlike the original GS method, which uses offline optimization for a specific scene to determine the 3D gaussian representation, we propose to train a neural network to directly generate an online 3D gaussian representation of the scene from a set of multi-view images. This representation is then rendered into a BeV feature map that is then analyzed by semantic segmentation heads.
The representation of a scene by a set of 3D gaussians allows to model its entire content. Indeed, the geometrical properties of the gaussians (position, size and rotation) enable to cover 3D space with varying level of detail depending on the structures encountered in the scene. Intuitively, a gaussian representing a lane marking will be rotated and elongated along its length. A gaussian representing a vehicle will be placed in the center of the vehicle and will follow its shape. The Figure 1(d) illustrates the intuition behind the representation proposed in this paper. Our contributions can be summarized as follows. (1) Introduction of GaussianBeV for BeV feature map generation from images through an online image-to-3D gaussian representation of the scene, allowing fine 3D content modeling. This representation is then splattered in BeV using a rasterizer module. To our knowledge, this is the first time that a gaussian splatting representation that is not scene-specific is proposed and integrated into a perception model. (2) Experiments demonstrating the effectiveness of our method, establishing it as the new state-of-the-art in BeV semantic segmentation.
@@ 2 Related work
Gaussian splatting. Gaussian splatting (GS) [12] is a 3D scene rendering technique which uses 3D gaussians to describe a scene. Each gaussian is parameterized by its position, scale, rotation, opacity and a Spherical Harmonics color model. The entire rendering pipeline is differentiable allowing the optimization of the gaussian parameters to a particular scene based on a set of images. GS is both fast and parallel, allowing real-time operation on GPUs. Moreover, it is possible to trade off details for rendering speed by pruning smaller and transparent gaussians. In comparison to sparse voxel grids, gaussians offer more efficient representations of a scene since individual gaussians can describe large volumes while smaller ones can accurately encode finer details with arbitrary resolution. Details and rasterization resolutions are parameters of the rendering pipeline, not the scene description. Several extensions have been proposed, allowing the management of dynamic objects [14, 25] or the distillation of semantic features from foundation models in the representation [20, 19, 29]. In all cases, the gaussian representation is scene-specific.
In our work, we propose to use gaussian representation to overcome the drawbacks of previous view transformer methods. Unlike previous gaussian splatting works, which learns gaussian representations offline, we propose to learn a neural network capable of providing an online gaussian representation of the scene.
@@ 3.1 Overview
The Figure 2 presents an overview of GaussianBeV. The model takes as input a set of multiview images $I\in\mathbb{R}^{N\times H\times W\times 3}$ with $N$ the number of cameras, $H$ and $W$ the dimensions of the images. These images are passed sequentially through four modules, leading to BeV segmentation.
@@
Given the input feature maps $F$, the 3D gaussian generator predicts the 3D gaussian representation of the scene using several prediction heads. The Figure 3 illustrates how it operates on the feature maps.
@@
Gaussian rotations. The 3D rotations of the gaussians in the scene are estimated by an rotation head applied to $F$. For a given pixel in the feature map of the camera $n$, it outputs an allocentric rotation in the form of a unit quaternion $q^{a}_{n,i}$. The allocentric rotation of a pixel corresponds to a rotation relative to the 3D optical ray passing through it. This modelization makes it easier for the rotation head to learn, as it has no knowledge of the optical ray corresponding to the pixel it is processing. To take an example, two objects placed at two different locations in the scene and with different absolute (egocentric) rotations in the camera reference frame may have the same appearance in the image. In this case, the allocentric rotation predicted by the rotation head will be the same. The intrinsic parameters of the camera are then used to retrieve the egocentric rotation information.
@@ 4.2 Ablations
(a) Ground truth
(b) -rot, -off
(c) +rot
(d) +rot, +off
Figure 6: Visualization of the 3D offsets and rotations predicted by the model by displaying their impact on the rendered BeV feature map in inference. We cropped region of interest for better visualization. (a) Ground truth, (b) disabling rotation and 3D offset prediction, (c) activating rotation prediction, (d) activating rotation and 3D offset prediction. Gaussians better fit the shape of objects when refined using these two predictions.
@@
Depth and early supervision. We investigate the influence of auxiliary losses on the gaussian representation as detailed in Sec 3.4. As shown in rows 1 and 2 of the Table 4, the use of the depth loss increases performance when GaussianBeV is trained without a BeV backbone. This shows that this loss regularizes the model to help it outputs a more geometrically consistent gaussian representation. Note that early supervision is used by default if the BeV backbone is not used, as this becomes the main segmentation loss. With regard to rows 3, 4 and 5 of the table using the entire BeV backbone, depth supervision and early supervision do not bring any performance gains when used separately. However, the last line of the table shows that, used together, these two losses improve results. In the Figure 5, we provide a visualization of the impact of early supervision on the 3D gaussian representation. We observe that using early supervision results in a more coherent BeV feature map. Indeed, the gaussians better fit the shape of the vehicles.
3D offsets and rotations. Table 4 (rows 6 to 8) shows the impact of learning 3D offsets and rotation. When GaussianBeV is trained without these two heads, it gives slightly poorer results than when it is learned with (last row). To visualize the 3D offsets and rotations learned by the model, these predictions are deactivated during inference to assess their impact on the BeV features. The Figure 6 shows that these predictions lead to a more accurate BeV map, more reliably fitting the shape of the objects.
@@ 5 Conclusion
In this paper, we introduced GaussianBeV, a novel image-to-BeV transformation method that is the new state-of-the-art on BeV semantic segmentation. Based on an online 3D gaussian generator, it transforms each pixel of an image feature map into semantized 3D gaussians. Gaussians are then splattered to obtain the BeV feature map. We have shown that the gaussian representation enables fine 3D modeling thanks to its ability to adapt to the different geometric structures present in the scene. We hope that this initial work will open the door to further research in 3D perception using online gaussian splatting representation.
6 Acknowledgements
This work benefited from the FactoryIA supercomputer financially supported by the Ile-de-France Regional Council.

### v2

@@ Abstract
The Bird’s-eye View (BeV) representation is widely used for 3D perception from multi-view camera images. It allows to merge features from different cameras into a common space, providing a unified representation of the 3D scene. The key component is the view transformer, which transforms image views into the BeV. However, actual view transformer methods based on geometry or cross-attention do not provide a sufficiently detailed representation of the scene, as they use a sub-sampling of the 3D space that is non-optimal for modeling the fine structures of the environment. In this paper, we propose GaussianBeV, a novel method for transforming image features to BeV by finely representing the scene using a set of 3D gaussians located and oriented in 3D space. This representation is then splattered to produce the BeV feature map by adapting recent advances in 3D representation rendering based on gaussian splatting [12]. GaussianBeV is the first approach to use this 3D gaussian modeling and 3D scene rendering process in an optimization free manner, i.e. without optimizing it on a specific scene and directly integrated into a single stage model for BeV scene understanding. Experiments show that the proposed representation is highly effective and place GaussianBeV as the new state-of-the-art on the BeV semantic segmentation task on the nuScenes dataset [2].
@@ 1 Introduction
In this article, we propose a novel view transformation method called GaussianBeV, which enables fine 3D modeling of scenes. Drawing on recent advances in explicit 3D representation for rendering novel views based on Gaussian Splatting (GS) [12], our method represents a scene using a set of 3D gaussians, each parameterized by a center, scale, rotation, opacity and semantic features (instead of colors in GS). Furthermore, unlike the original GS method, which uses offline optimization for a specific scene to determine the 3D gaussian representation, we propose to train a neural network to directly generate an optimization free 3D gaussian representation of the scene from multi-view images. This representation is then rendered into a BeV feature map which is analyzed by semantic segmentation heads.
The representation of a scene by a set of 3D gaussians allows to model its entire content. Indeed, the geometrical properties of the gaussians (position, size and rotation) enable to cover 3D space with varying level of detail depending on the structures encountered in the scene. Intuitively, a gaussian representing a lane marking will be rotated and elongated along its length. A gaussian representing a vehicle will be placed in the center of the vehicle and will follow its shape. Figure 1(d) illustrates the intuition behind the representation proposed in this paper. Our contributions can be summarized as follows. (1) Introduction of GaussianBeV for BeV feature map generation from images through an optimization free image-to-3D gaussian representation for any scene, allowing fine 3D content modeling. This representation is then splattered in BeV using a rasterizer module. To our knowledge, this is the first time that a gaussian splatting representation that is not scene-specific is proposed and integrated into a BeV perception model. (2) Experiments demonstrating the effectiveness of our method, establishing it as the new state-of-the-art in BeV semantic segmentation.
@@ 2 Related work
Gaussian splatting. Gaussian splatting (GS) [12] is a 3D scene rendering technique which uses 3D gaussians to describe a scene. Each gaussian is parameterized by its position, scale, rotation, opacity and a Spherical Harmonics color model. The entire rendering pipeline is differentiable allowing the optimization of the gaussian parameters to a particular scene based on a set of images. GS is both fast and parallel, allowing real-time operation on GPUs.
In comparison to sparse voxel grids, gaussians offer more efficient representations of a scene since individual gaussians can describe large volumes while smaller ones can accurately encode finer details with arbitrary resolution. Several extensions have been proposed, allowing the management of dynamic objects [14, 28] or the distillation of semantic features from foundation models in the representation [22, 21, 34, 30]. Although GS rasterization itself is fast, the optimization of gaussian parameters to match a scene usually requires many iterations. In order to address this issue for real-time SLAM, [18] leverages ground-truth 3D information from a depth sensor to initialize gaussian positions and [27] uses temporal consistency in order to minimize the optimization workload between successive frames. To perform scene reconstruction with stereo images [4] or object-centric single image [23], optimization-free methods are introduced allowing the model to directly learn to predict the gaussian parameters.
In our work, we propose to overcome the drawbacks of view transformer methods by modelling scenes as set of gaussians which are subsequently rasterized into a feature map. Unlike previous gaussian splatting works which optimize the gaussian representations to each scene, we propose to learn a neural network capable of directly predicting a gaussian representation of any scene. Compared to [4, 23] which targets scene reconstruction, our model is designed to perform real-time outdoor BeV semantic segmentation from multiple calibrated cameras.
@@ 3.1 Overview
Figure 2 presents an overview of GaussianBeV. The model takes as input a set of multiview images $I\in\mathbb{R}^{N\times H\times W\times 3}$ with $N$ the number of cameras, $H$ and $W$ the dimensions of the images. These images are passed sequentially through four modules, leading to BeV segmentation.
@@
Given the input feature maps $F$, the 3D gaussian generator predicts the 3D gaussian representation of the scene using several prediction heads. Figure 3 illustrates how it operates on the feature maps.
@@
Gaussian rotations. The 3D rotations of the gaussians in the scene are estimated by a rotation head applied to $F$. For a given pixel in the feature map of the camera $n$, it outputs an allocentric rotation in the form of a unit quaternion $q^{a}_{n,i}$. The allocentric rotation of a pixel corresponds to a rotation relative to the 3D optical ray passing through it. This modelization makes it easier for the rotation head to learn, as it has no knowledge of the optical ray corresponding to the pixel it is processing. To take an example, two objects placed at two different locations in the scene and with different absolute (egocentric) rotations in the camera reference frame may have the same appearance in the image. In this case, the allocentric rotation predicted by the rotation head will be the same. The intrinsic parameters of the camera are then used to retrieve the egocentric rotation information.
@@ 4.2 Ablations
Depth and early supervision. We investigate the influence of auxiliary losses on the gaussian representation as detailed in Sec 3.4. As shown in rows 1 and 2 of the Table 4, the use of the depth loss increases performance when GaussianBeV is trained without a BeV backbone. This shows that this loss regularizes the model to help it outputs a more geometrically consistent gaussian representation. Note that early supervision is used by default if the BeV backbone is not used, as this becomes the main segmentation loss. With regard to rows 3, 4 and 5 of the table using the entire BeV backbone, depth supervision and early supervision do not bring any performance gains when used separately. However, the last line of the table shows that, used together, these two losses improve results. In Figure 5, we provide a visualization of the impact of early supervision on the 3D gaussian representation. We observe that using early supervision results in a more coherent BeV feature map. Indeed, the gaussians better fit the shape of the vehicles.
3D offsets and rotations. Table 4 (rows 6 to 8) shows the impact of learning 3D offsets and rotation. When GaussianBeV is trained without these two heads, it gives slightly poorer results than when it is learned with (last row).
@@ 5 Conclusion
In this paper, we introduced GaussianBeV, a novel image-to-BeV transformation method that is the new state-of-the-art on BeV semantic segmentation. Based on an optimization free 3D gaussian generator, it transforms each pixel of an image feature map into semantized 3D gaussians. Gaussians are then splattered to obtain the BeV feature map. We have shown that the gaussian representation enables fine 3D modeling thanks to its ability to adapt to the different geometric structures present in the scene. Future work will aim to generate a more compact 3D Gaussian representation to improve computing time. We hope that this initial work will open the door to further research in 3D perception using optimization free gaussian splatting representation.
Acknowledgements. This work was carried out as part of the CEA-Valeo joint laboratory and partially financed by BPI France as part of the AVFS project. This work benefited from the FactoryIA supercomputer financially supported by the Ile-de-France Regional Council.